OpenStudy (anonymous):

A large emerald with a mass of 812.04 grams was recently discovered in a mine. If the density of the emerald is 4.02 grams/cm^2 , what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

@ganeshie8 can you help me please

ganeshie8 (ganeshie8):

\(\large Volume = \frac{Mass}{density}\)

OpenStudy (anonymous):

i know the mass is 812.04 grams but what is the density?

ganeshie8 (ganeshie8):

\(\large Volume = \frac{812.04 grams}{4.02 grams/cm^\color{red}{3}}\)

OpenStudy (anonymous):

can just put that in my calculator

ganeshie8 (ganeshie8):

\(\large Volume = \frac{812.04 }{4.02}~cm^\color{red}{3}\)

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

ok im about to

OpenStudy (anonymous):

202 cm^3 for future people!